OpenStudy (anonymous):

Solve each equation. Show all work 1.(1/6x2 ) = ( 1/3x2) - (1/x) 2. (1/n2) + (1/n) = (1/2n2) 3. (1/x) = (6/3x) + 1 4. (1/6x2) = (1/2x) + (7/ 6x2) 5. (1/6b2) + (1/6b) = (1/b2)

OpenStudy (anonymous):

help me anyone

OpenStudy (anonymous):

help help anyone PLEASE

OpenStudy (mathmale):

Please choose ONE question, and then we could focus on that one. Discuss what you already know and what you need to know to solve this question.

OpenStudy (anonymous):

1.(1/6x2 ) = ( 1/3x2) - (1/x)

OpenStudy (tkhunny):

Is it supposed to be this? \(\dfrac{1}{6x^{2}} = \dfrac{1}{3x^{2}} - \dfrac{1}{x}\) In any case, you would do well to find a common denominator and add things up. Go!!

OpenStudy (anonymous):

yes now wat i supose to do

OpenStudy (mathmale):

As tkhunny has aptly pointed out, you need to determine what the common denominator is in this particular problem.

OpenStudy (anonymous):

aint the common denominator is 6

OpenStudy (mathmale):

Hint; IF WE were to add 1/3 and 1/7, we'd identify the lowest common denominator to be 3*7 = 21, and then change each fraction accordingly: 1/3 = 7/21 1/7 = 3/21 1/3 + 1/7 = (3+7)/21 = 10/21. Think: how could you use this information / example to find the LCD of your posted problem?

OpenStudy (anonymous):

6*3=18 1/3=6/18 1/6=3/18 1/3 + 1/6=(3+6)/18=9/18

OpenStudy (mathmale):

Please remember that what I typed in most recently was an example. You'll need to apply the same principle(s) towards solving (1/6x2 ) = ( 1/3x2) - (1/x). \[(1/6x^2 ) = ( 1/3x^2) - (1/x)\] hs common denominator 6x^2. Note that both 3x^2 and x will divide into 6x^2. \[(1/6x^2 ) = ( 1/3x^2) - (1/x) \rightarrow \frac{ 1 }{ 6x^2 }=\frac{ 1 }{ 3x^2 }-\frac{ 1 }{ x }\]

OpenStudy (mathmale):

can be re-written with the common denominator in all terms as:\[\frac{ 1 }{ 6x^2 }=\frac{ 1(2) }{ 2(3x^2) }-\frac{ 1(6x) }{ x(6x) }\]

OpenStudy (mathmale):

and then, because the denoms. are now all the same, you can ignore them and concentrate on solving 1=2-6x. I need to get off OpenStudy now. But I hope to be able to work with you again soon.